Ruth is a senior citizen who consumes food and health care. When she retired in​ 2002, she consumed 90 meals and 3 medical checkups per month. Suppose in 2002 that the price of a meal was ​$10 and the price of a medical checkup was ​$57. ​However, in​ 2012, the price of a meal is ​$11 and the price of a medical checkup is ​$306. As a​ result, Ruth consumes 100 meals and 2 medical checkups per month. Calculate a Laspeyres​ cost-of-living index for Ruth using 100 as the base in 2002. In​ particular, the Laspeyres index for 2012 is nothing. ​(Enter your response rounded to two decimal places.​)
Answer #1

Laspeyres Price Index in Year 2012 = Price of individual item in the observation period * Quantity of Individual item in base period / Price of individual item in base period * Quantity of individual item in base period * 100

Laspeyres Price Index = 11*90 + 306*3 / 10*90 + 57*3 *100

= 990 + 918 / 900 + 171 *100

= 1908 / 1071 * 100

= 1.782 * 100

= 178.2